# MUSE spectroscopy of the high abundance discrepancy planetary nebula NGC 6153

### v2

@@ 4.3 Correction for recombination contribution
In the case of the auroral lines [N ii]5755 and [O ii]7320+30, it is well known that they are also produced by recombination, an emission process favored at low temperatures. In the bulk of H ii regions and PNe, this contribution is relatively small and the correction is much smaller than the typical uncertainties. However, this is not the case for high-ADF PNe, where extremely strong recombination emission of both N ii and O ii lines from an hypothetical cold emission plasma can significantly contribute to the observed flux of these lines (see Gómez-Llanos et al., 2020). Following the procedure described in Sec. 5.2 of GR22, we correct the observed emission of [N ii]5755 and [O ii]7320+30 from the contribution due to recombination using N ii5679 and O ii4649+50 respectively by applying the following relations:
@@
![Refer to caption](x5.png)
![Refer to caption](x6.png)
![Refer to caption](x7.png)
@@ 4.4 Physical conditions
S2+ | Froese Fischer et al. (2006) | Tayal & Gupta (1999a) |
@@ 4.5 Ionic chemical abundance maps
Ionic chemical abundance maps are constructed only for the emission lines with the highest signal-to-noise ratio for each ion detected in the NGC 6153 spectrum. To determine the ionic abundances from these emission lines, we have to define the and values to be used in computing the line emissivities and the H emissivity. The case of He+/H+ and He2+/H+ is relatively complicated, as there is no single plasma component emitting helium, but rather two different components: warm and cold. We will discuss them separately in Sect. A. For the RLs of heavy elements, we use the and values of 2,000 K and 10,000 cm-3 obtained by R22. For the CELs, we use the values derived from the [N ii]5755/6548[S ii]6716/6731 for IP¡17 eV, and [S iii]6312/9069[Cl iii]5518/5538 for IP17 eV. Owing to the noisy maps of ([Ar iii]) and ([Ar iv]) (see Fig. 8) we did not make the attempt to use them to compute ionic abundances; however in Sect. 5.2 we will discuss the effect on the ionic abundances of adopting these temperatures for high ionization potential ions (IP 35 eV).
@@
For the H line, the usual methodology is to adopt and selected for the specific emission line (RL or CEL) in question, following the “classical” approach based on eq. 3 from Morisset et al. (2023). A second order effect can be taken into account using a specific value for corresponding to the H i emission lines, as discussed by Morisset et al. (2023) at the end of their Sec. 2. In the present case, for spatially resolved maps, this value of (H i) would fall between ([N ii]) and ([S iii]). Since these two temperatures are very close, we adhere to the classical method where (H i) is the same as the ionic temperature of the considered ion. Regarding the integrated spectrum, we discuss the effects of using an ad-hoc value for (H i) in Sec. 5.2.
![Refer to caption](x10.png)
@@
The abundance map of O ii4649.13 + 4650.85 is of particular interest as it is the one used to compute the ADF maps. However, considering the relatively low spectral resolution of the MUSE data and the results of Liu et al. (2000) and McNabb et al. (2016), these lines can be blended with C iii4650.25. Liu et al. (2000) reported that C iii4650.25 contributed 9 per cent to the sum of O ii4649.13 + 4650.85 and C iii4650.25 integrated fluxes. In Fig. 10 we show the oxygen abundance maps derived from O ii4649.13 + 4650.85 (left panel), and the lower signal-to-noise map from O ii4661.63 line (central panel). In the right panel of Fig. 10 we show the histograms of the O2+/H+ abundance ratio obtained from the O ii4649.13 + 4650.85 blend (blue) and from the O ii4661.63 RL (orange). It is clear that both histograms peak at very similar values but show slightly different distributions. The long tail of O ii4661.63 corresponds to the very high values found in the most external zones of the map, where the signal-to-noise of the line is lower. However, the differences in the distributions, evident in the left wings of both histograms, are the result of different number of spaxels with detected line emissions in each scenario. In this context, no potential contamination of the C iii4650.25 line can be inferred from these histograms, given the close proximity of their peaks.
@@ 4.6 Weight of the cold component in the H emission
As Bohigas (2015) and Morisset et al. (2023) have pointed out, for PNe with two plasma components with significantly different physical conditions, the calculation of chemical abundances must take into account the weight of the line emission in both the cold and warm components, rather than assuming homogeneous physical conditions. We have followed the methodology described by Morisset et al. (2023) to determine the chemical abundances, considering the weight of the cold region in the H emission (, see eqs. 5, 6, 7 and 8 in Morisset et al. 2023). It is worth mentioning that for the case of He ions, the case is much more complicated since ionic He lines may be emitted by both gas phases if the cold region is not He-poor. In Sect. A we discuss such a case.
@@
The normalized Paschen jump PJ/I9229 is obtained for a grid of (from 6,000 to 10,000 K), (from 1,800 to 3,000 K), and (from 0.0 to 0.15), and constant values of = 3,000 cm-3, = 10cm-3, He+/H+ = 0.1, and He2+/H+ = 0.005 (Richer et al., 2022).
@@
![Refer to caption](x12.png)
Once we have the maps of and the ionic abundances, we generate the maps of the classical ADF, obtained from the ratio of abundance obtained from the RLs and CELs, computed using the abundance values without taking into account . We also generate maps of the same abundance ratios, but taking into account the presence of the cold region and applying : this is what we call the Abundance Contrast Factor i.e. ACF (see Morisset et al., 2023). These maps are shown in Fig. 13, where the left (right) panels show the ADF (ACF) for O+/H+ (O2+/H+) in the upper (lower) panels. The effect of the cold region is clear in enhancing the contrast between both abundances by an average of 0.9 dex. At this point, we have to emphasize that the spatial distribution of the ADF maps is equivalent to the corresponding RL/CEL line ratios. This is an important point because the different spatial distributions of the emissivities of RLs and CELs for a given ion are not evident from Fig. 16, contrary to the objects presented in GR22 (see their Figs. S1 to S3).
![Refer to caption](x13.png)
@@ 5 The integrated emission line spectrum of NGC 6153
To compute elemental abundances of elements heavier than helium we have to take into account unseen ionisation stages of the different elements by adopting ionisation correction factors (ICFs). GR22 pointed out that computing ICFs only makes sense when dealing with integrated spectra of the whole PN or, at least, spectra covering a significant volume of the PN that roughly includes the whole ionisation structure of the nebula. This is not the case of single spaxels of IFU observations so elemental abundances are only computed from the integrated spectrum within the mask described in Sect. 3. We then apply the extinction correction following the same methodology presented in Sect. 4.2 to obtain the extinction-corrected line fluxes. Unreddened and extinction-corrected line fluxes along with their uncertainties are presented in Table C. The reported extinction-corrected line fluxes of [N ii]5755 and [O ii]7320+30 were corrected from recombination contribution assuming = 6,000 K and = 10,000 cm-3 (see Sect. 4.3). Lines affected by blends are also indicated in the table.
We used Monte Carlo simulations to estimate the uncertainties in the physical conditions and chemical abundances from the integrated spectrum. We generated 500 random values for each line intensity using a Gaussian distribution centred on the observed line intensity with a standard deviation equal to its uncertainty, then the extinction coefficient, physical conditions and chemical abundances are determined for the 501 values of line intensities, and their corresponding uncertainties are determined using standard deviation or 16-84 percentiles.
@@ 5.1 Physical conditions
We computed the physical conditions ( and ) from the integrated spectrum using several pairs of diagnostics. As mentioned in Sect. 4.3, and following the same procedure, we have considered a 40 per cent recombination contribution to the auroral [N ii]5755 line for the computation of ([N ii]), leading to a corrected of 8760 K, i.e, 1800 K lower than without taking into account such a correction. Similarly, we also applied this procedure to the intensities of the [O ii]7320+30 which leads to a 51 per cent recombination contribution. In Table C, we present for these lines both the unreddened measured fluxes as well as the intensities corrected for extinction and recombination contribution.
As is well known, the main drawback of MUSE spectra of photoionised nebulae is the lack of coverage of wavelengths bluer than 4600 Å preventing the detection of the auroral [O iii] 4363 line, which is essential for computing the widely used ([O iii]). However, the depth of the MUSE NGC 6153 spectrum allows the detection of additional -sensitive auroral lines in the integrated spectrum with enough signal-to-noise to compute electron temperatures. This is exemplified by the faint [Ar iii] 5192 and [Ar iv] 7170 lines. The calculated physical conditions are presented in Table 5.1. To compute (PJ) and (He i) we assumed the density provided by the ( ([S iii]), ([S ii]) pair. We also computed an averaged value of ([Ar iii]) and ([Ar iv]), given by:
@@
The electron temperature derived from the [S iii] line ratio is lower than that obtained from other ions. We have used the PyNeb default set of atomic data labeled PYNEB_23_01. We made the test of modifying the default [S iii] effective collision strengths from Tayal & Gupta (1999b) to Grieve et al. (2014), and obtained a ([S iii]) 300 K higher, which would be in better agreement with what is obtained from the [Ar iii] lines. The effect of these new atomic data in the ionic abundance maps shown in Fig. 12 would be relatively small and the resulting spatial distributions would be practically unaffected.
@@ 5.2 Ionic abundances
a ([N ii]) for ions with IP ¡ 17 eV / ([S iii]) for ions with IP ¿ 17 eV
b “Ion” meaning that for computing the emissivity of H i we adopt the same physical conditions than for the corresponding ion; “PJ” is the computed from the Paschen jump.
c For the warm (w) and cold (c) regions. See Sect. 5.3.
@@
Our favored recipe is the number 7, where each ion emissivity (including H+) is computed using its corresponding physical conditions, and the factor is taken into account.
@@ 5.3 The helium case
The determination of He/H for both regions is then: 0.089 ¡ (He/H)w ¡ 0.10, and 0.30 ¡ (He/H)c ¡ 0.38. The overabundance of helium in the cold region is much smaller than that of the metals. This result confirms the upper limit of He/H ¡ 6 in the cold region obtained by Gómez-Llanos & Morisset (2020). The emission of the dominant ion He+ that comes from each region are comparable: 60 per cent (70 per cent) of the He i 6678 (7281, respectively) comes from the warm phase.
@@ 5.4 Total abundances
To calculate the elemental abundances of elements heavier than helium in the warm phase, an ionisation correction factor (ICF) should be computed to account for unobserved ionisation states. Similarly to GR22, in this work we have taken three different approaches to calculate ICFs: i) using ICFs from the literature; ii) looking for ICFs from photoionisation models close to the observed object, and iii) computing ad hoc ICFs using machine learning techniques.
@@
database (Morisset et al., 2015) to train an Artificial Neural Network (ANN) to predict the needed ICFs from some ionic abundance ratios. In this case, we select models that fit the observed values of the O2+/O+, S2+/S+, Cl3+/Cl2+ and Ar3+/Ar2+ ionic abundance ratios within 0.5 dex, and He2+/He+ abundance ratio within 0.1 dex. We use 80 per cent of this set of models to train the ANN (16,756 models) and test the prediction of the ANN on the remaining 20 per cent (4,189 models). The input vector is built from the 5 ionic abundance ratios described above. The output vector contains 8 ICFs: ICF(N+/H+), ICF((O+ + O2+)/H+), ICF((S+ + S2+)/H+), ICF((Cl2+ + Cl3+)/H+), ICF((Ar2+ + Ar3+)/H+), ICF(N+/O+), ICF((S+ + S2+)/O+), and ICF(Ar2+/(O+ + O2+)). The ANN is obtained using the SciKit-Learn Python library. It contains 3 dense layers of 50, 80, and 80 RELU perceptrons, respectively. Convergence is reached using the ADAM algorithm.
@@
![Refer to caption](x14.png)
![Refer to caption](x15.png)
![Refer to caption](x16.png)
![Refer to caption](x17.png)
![Refer to caption](x18.png)
![Refer to caption](x19.png)
![Refer to caption](x20.png)
![Refer to caption](x21.png)
![Refer to caption](x22.png)
![Refer to caption](x23.png)
![Refer to caption](x24.png)
@@
a In their Tab. 3, Delgado-Inglada et al. (2014) wrongly parameterize their ICFf using while it should be using , as they correctly write in their text.
@@
In Table 7, we present the elemental abundances obtained from the three methods applied in this work. A comparison between the ICFs obtained using different techniques results in elemental abundances that are in relatively good agreement for all cases. It is worth mentioning that although the He2+/He fraction in each plasma component of the nebula is quite uncertain, its low value has almost no effect on the only ICF we have used from Delgado-Inglada et al. (2014), which is parameterized using this ratio (i.e., ICF(O+ + O2+)), minimising possible systematic effects on our determinations. In Fig. 14 we show the Monte Carlo distributions obtained for the elemental abundances using the ICFs shown in Table 6; only the ICF distributions obtained using more than one method are shown. The different approaches for computing the ICFs (literature, look-up table and ANN are shown in purple, green, and orange, respectively). The vertical lines on each colour show the median of the distribution.

### v3

@@ 4.3 Correction for recombination contribution
In the case of the auroral lines [N ii]5755 and [O ii]7320+30, it is well known that they are also produced by recombination, an emission process favoured at low temperatures. In the bulk of H ii regions and PNe, this contribution is relatively small, and the correction is much smaller than the typical uncertainties. However, this is not the case for high-ADF PNe, where extremely strong recombination emission of both N ii and O ii lines from an hypothetical cold emission plasma can significantly contribute to the observed flux of these lines (see Gómez-Llanos et al., 2020). Following the procedure described in Sec. 5.2 of GR22, we correct the observed emission of [N ii]5755 and [O ii]7320+30 from the contribution due to recombination using N ii5679 and O ii4649+50 respectively by applying the following relations:
@@ 4.4 Physical conditions
S2+ | Froese Fischer et al. (2006) | Tayal & Gupta (1999) |
@@ 4.5 Ionic chemical abundance maps
Ionic chemical abundance maps are constructed only for the emission lines with the highest signal-to-noise ratio for each ion detected in the NGC 6153 spectrum. To determine the ionic abundances from these emission lines, we have to define the and values to be used in computing the line emissivities and the H emissivity. The case of He+/H+ and He2+/H+ is relatively complicated, as there is no single plasma component emitting helium, but rather two different components: warm and cold. We will discuss them separately in Sect. A. For the RLs of heavy elements, we use the and values of 2,000 K and 10,000 cm-3 obtained by R22. For the CELs, we use the values derived from the [N ii]5755/6548[S ii]6716/6731 for IP¡17 eV, and [S iii]6312/9069[Cl iii]5518/5538 for IP17 eV. Owing to the noisy maps of ([Ar iii]) and ([Ar iv]) (see Fig. 8) we did not attempt to use them to compute ionic abundances; however, in Sect. 5.2 we will discuss the effect on the ionic abundances of adopting these temperatures for high ionisation potential ions (IP 35 eV).
@@
For the H line, the usual methodology is to adopt and selected for the specific emission line (RL or CEL) in question, following the “classical” approach based on eq. 3 from Morisset et al. (2023). A second order effect can be taken into account using a specific value for corresponding to the H i emission lines, as discussed by Morisset et al. (2023) at the end of their Sec. 2. In the present case, for spatially resolved maps, this value of (H i) would fall between ([N ii]) and ([S iii]). Since these two temperatures are very close, we adhere to the classical method, where (H i) is the same as the ionic temperature of the considered ion. Regarding the integrated spectrum, we discuss the effects of using an ad-hoc value for (H i) in Sec. 5.2.
@@
The abundance map of O ii4649.13 + 4650.85 is of particular interest as it is the one used to compute the ADF maps. However, considering the relatively low spectral resolution of the MUSE data and the results of Liu et al. (2000) and McNabb et al. (2016), these lines can be blended with C iii4650.25. Liu et al. (2000) reported that C iii4650.25 contributed 9 per cent to the sum of O ii4649.13 + 4650.85 and C iii4650.25 integrated fluxes. In Fig. 10 we show the oxygen abundance maps derived from O ii4649.13 + 4650.85 (left panel), and the lower signal-to-noise map from O ii4661.63 line (central panel). In the right panel of Fig. 10 we show the histograms of the O2+/H+ abundance ratio obtained from the O ii4649.13 + 4650.85 blend (blue) and from the O ii4661.63 RL (orange). It is clear that both histograms peak at very similar values but show slightly different distributions. The long tail of O ii4661.63 corresponds to the very high values found in the most external zones of the map, where the signal-to-noise of the line is lower. However, the differences in the distributions, evident in the left wings of both histograms, are the result of different number of spaxels with detected line emissions in each scenario. In this context, no potential contamination of the C iii4650.25 line can be inferred from these histograms, given the proximity of their peaks.
@@ 4.6 Weight of the cold component in the H emission
As Bohigas (2015) and Morisset et al. (2023) have pointed out, for PNe with two plasma components with significantly different physical conditions, the calculation of chemical abundances must take into account the weight of the line emission in both the cold and warm components, rather than assuming homogeneous physical conditions. We have followed the methodology described by Morisset et al. (2023) to determine the chemical abundances, considering the weight of the cold region in the H emission (, see eqs. 5, 6, 7 and 8 in Morisset et al. 2023). It is worth mentioning that for the case of He ions, the case is much more complicated, since ionic He lines may be emitted by both gas phases if the cold region is not He-poor. In Sect. A we discuss such a case.
@@
The normalised Paschen jump PJ/I9229 is obtained for a grid of (from 6,000 to 10,000 K), (from 1,800 to 3,000 K), and (from 0.0 to 0.15), and constant values of = 3,000 cm-3, = 10cm-3, He+/H+ = 0.1, and He2+/H+ = 0.005 (Richer et al., 2022).
@@
Once we have the maps of and the ionic abundances, we generate the maps of the classical ADF, obtained from the ratio of abundance obtained from the RLs and CELs, computed using the abundance values without taking into account . We also generate maps of the same abundance ratios, but taking into account the presence of the cold region and applying : this is what we call the abundance contrast factor, that is, ACF (see Morisset et al., 2023). These maps are shown in Fig. 13, where the left (right) panels show the ADF (ACF) for O+/H+ (O2+/H+) in the upper (lower) panels. The effect of the cold region is clear in enhancing the contrast between both abundances by an average of 0.9 dex. At this point, we have to emphasize that the spatial distribution of the ADF maps is equivalent to the corresponding RL/CEL line ratios. This is an important point because the different spatial distributions of the emissivities of RLs and CELs for a given ion are not evident from Fig. 16, contrary to the objects presented in GR22 (see their Figs. S1 to S3).
@@ 5 The integrated emission line spectrum of NGC 6153
To compute elemental abundances of elements heavier than helium, we have to take into account unseen ionisation stages of the different elements by adopting ionisation correction factors (ICFs). GR22 pointed out that computing ICFs only makes sense when dealing with integrated spectra of the whole PN or, at least, spectra covering a significant volume of the PN that roughly includes the whole ionisation structure of the nebula. This is not the case of single spaxels of IFU observations, so elemental abundances are only computed from the integrated spectrum within the mask described in Sect. 3. We then apply the extinction correction following the same methodology presented in Sect. 4.2 to obtain the extinction-corrected line fluxes. Unreddened and extinction-corrected line fluxes along with their uncertainties are presented in Table 9. The reported extinction-corrected line fluxes of [N ii]5755 and [O ii]7320+30 were corrected from recombination contribution assuming = 6,000 K and = 10,000 cm-3 (see Sect. 4.3). Lines affected by blends are also indicated in the table.
We used Monte Carlo simulations to estimate the uncertainties in the physical conditions and chemical abundances from the integrated spectrum. We generated 500 random values for each line intensity using a Gaussian distribution centred on the observed line intensity with a standard deviation equal to its uncertainty, then the extinction coefficient, physical conditions and chemical abundances are determined for the 501 values of line intensities, and their corresponding uncertainties are determined using standard deviation or 16–84 percentiles.
@@ 5.1 Physical conditions
We computed the physical conditions ( and ) from the integrated spectrum using several pairs of diagnostics. As mentioned in Sect. 4.3, and following the same procedure, we have considered a 40 per cent recombination contribution to the auroral [N ii]5755 line for the computation of ([N ii]), leading to a corrected of 8760 K, i.e, 1800 K lower than without taking into account such a correction. Similarly, we also applied this procedure to the intensities of the [O ii]7320+30 which leads to a 51 per cent recombination contribution. In Table 9, we present for these lines both the unreddened measured fluxes and the intensities corrected for extinction and recombination contribution.
As is well known, the main drawback of MUSE spectra of photoionised nebulae is the lack of coverage of wavelengths bluer than 4600 Å preventing the detection of the auroral [O iii] 4363 line, which is essential for computing the widely used ([O iii]). However, the depth of the MUSE NGC 6153 spectrum allows the detection of additional -sensitive auroral lines in the integrated spectrum with enough signal-to-noise to compute electron temperatures. This is exemplified by the faint [Ar iii] 5192 and [Ar iv] 7170 lines. The calculated physical conditions are presented in Table 4. To compute (PJ) and (He i) we assumed the density provided by the ( ([S iii]), ([S ii]) pair. We also computed an averaged value of ([Ar iii]) and ([Ar iv]), given by:
@@
The electron temperature derived from the [S iii] line ratio is lower than that obtained from other ions. We have used the PyNeb default set of atomic data labelled PYNEB_23_01. We made the test of modifying the default [S iii] effective collision strengths from Tayal & Gupta (1999) to Grieve et al. (2014), and obtained a ([S iii]) 300 K higher, which would be in better agreement with what is obtained from the [Ar iii] lines. The effect of these new atomic data in the ionic abundance maps shown in Fig. 12 would be relatively small, and the resulting spatial distributions would be practically unaffected.
@@ 5.2 Ionic abundances
a The first 6 rows summarise the temperatures used in each recipe. The following rows give the ionic abundance obtained from the emission line given in the corresponding first column.
b ([N ii]) for ions with IP ¡ 17 eV / ([S iii]) for ions with IP ¿ 17 eV.
c “Ion” meaning that for computing the emissivity of H i we adopt the same physical conditions than for the corresponding ion; “PJ” is the computed from the Paschen jump.
d For the warm (w) and cold (c) regions. See Sect. 5.3.
@@
Our favoured recipe is the number 7, where each ion emissivity (including H+) is computed using its corresponding physical conditions, and the factor is taken into account.
@@ 5.3 The helium case
The determination of He/H for both regions is then: 0.089 ¡ (He/H)w ¡ 0.10, and 0.30 ¡ (He/H)c ¡ 0.38. The overabundance of helium in the cold region is much smaller than that of the metals. This result confirms the upper limit of He/H ¡ 6 in the cold region, obtained by Gómez-Llanos & Morisset (2020). The emission of the dominant ion He+ that comes from each region are comparable: 60 per cent (70 per cent) of the He i 6678 (7281, respectively) comes from the warm phase.
@@ 5.4 Total abundances
To calculate the elemental abundances of elements heavier than helium in the warm phase, an ICF should be computed to account for unobserved ionisation states. Similarly to GR22, in this work we have taken three different approaches to calculate ICFs: i) using ICFs from the literature; ii) looking for ICFs from photoionisation models close to the observed object, and iii) computing ad hoc ICFs using machine learning techniques.
@@
database (Morisset et al., 2015) to train an Artificial Neural Network (ANN) to predict the needed ICFs from some ionic abundance ratios. In this case, we select models that fit the observed values of the O2+/O+, S2+/S+, Cl3+/Cl2+ and Ar3+/Ar2+ ionic abundance ratios within 0.5 dex, and He2+/He+ abundance ratio within 0.1 dex. We use 80 per cent of this set of models to train the ANN (16,756 models) and test the prediction of the ANN on the remaining 20 per cent (4,189 models). The input vector is built from the five ionic abundance ratios described above. The output vector contains eight ICFs: ICF(N+/H+), ICF((O+ + O2+)/H+), ICF((S+ + S2+)/H+), ICF((Cl2+ + Cl3+)/H+), ICF((Ar2+ + Ar3+)/H+), ICF(N+/O+), ICF((S+ + S2+)/O+), and ICF(Ar2+/(O+ + O2+)). The ANN is obtained using the SciKit-Learn Python library. It contains 3 dense layers of 50, 80, and 80 RELU perceptrons, respectively. Convergence is reached using the ADAM algorithm.
@@
a In their Tab. 3, Delgado-Inglada et al. (2014) wrongly parameterise their ICFf using while it should be using , as they correctly write in their text.
@@
In Table 7, we present the elemental abundances obtained from the three methods applied in this work. A comparison between the ICFs obtained using different techniques results in elemental abundances that are in relatively good agreement for all cases. It is worth mentioning that although the He2+/He fraction in each plasma component of the nebula is quite uncertain, its low value has almost no effect on the only ICF we have used from Delgado-Inglada et al. (2014), which is parameterised using this ratio (i.e. ICF(O+ + O2+)), minimising possible systematic effects on our determinations. In Fig. 14 we show the Monte Carlo distributions obtained for the elemental abundances using the ICFs shown in Table 6; only the ICF distributions obtained using more than one method are shown. The different approaches for computing the ICFs (literature, look-up table and ANN are shown in purple, green, and orange, respectively). The vertical lines on each colour show the median of the distribution.